OpenStudy (crashonce):

The sum of an infinite GP is 6. The Sum of the first two terms is 4.5. What is the first term

OpenStudy (crashonce):

I've done this but my answer is wrong

OpenStudy (crashonce):

because my r was root -2

OpenStudy (anonymous):

Hmmm, my guess is: \[ a+ar = 4.5\\ 6=\frac{a}{1-r} \]

OpenStudy (crashonce):

thats what i got

OpenStudy (crashonce):

what is your answer

OpenStudy (anonymous):

First I do:\[ r = 1-\frac{a}{6} \]

OpenStudy (anonymous):

then\[ a\left(1+\left(1-\frac a6\right)\right) = 4.5 \]

OpenStudy (anonymous):

it's a quadratic equation I guess.

OpenStudy (crashonce):

so 9 or 3

OpenStudy (anonymous):

sum to infinity is a/1-r => first term and second terms are a,ar a(1+r) = 4.5 - eq1 a/1-r = 6 - eq2 eq1/eq2 1-r2 = 4.5/6 = 3/4 r2 = 1/4 r = 1/2 or -1/2 a(3/2) = 6 a =4 or 12 Hope its Correct...

OpenStudy (anonymous):

Ahahaha.

OpenStudy (anonymous):

Isn't That Correct? :\

OpenStudy (crashonce):

not a choice for the answers